# Scalable Oversight for Superhuman AI via Recursive Self-Critiquing

### v3

@@ Abstract
As AI capabilities increasingly surpass human proficiency in complex tasks, current alignment techniques including SFT and RLHF face fundamental challenges in ensuring reliable oversight. These methods rely on direct human assessment and become untenable when AI outputs exceed human cognitive thresholds. In response to this challenge, we explore two hypotheses: (1) Critique of critique can be easier than critique itself, extending the widely-accepted observation that verification is easier than generation to the critique domain, as critique itself is a specialized form of generation; (2) This difficulty relationship is recursively held, suggesting that when direct evaluation is infeasible, performing high-order critiques (e.g., critique of critique of critique) offers a more tractable supervision pathway. We further conduct Human-AI and AI-AI experiments to investigate the potential of utilizing recursive self-critiquing for AI supervision. Our results highlight recursive critique as a promising approach for scalable AI oversight.
@@ 1 Introduction
The provision of supervision signals is fundamental to AI alignment (Bowman et al., 2022). From the supervision signal acquisition perspective, tasks can be categorized as: (1) tasks with well-defined criteria, where ground truth can be deterministically obtained with low computational overhead, e.g., Go games and mathematical problems (Silver et al., 2017; Lightman et al., 2023); (2) tasks with subjectivity or complex evaluation frameworks, such as business strategy and product design jobs (Ouyang et al., 2022). Tasks of the second type are more prevalent in real-world applications and predominantly rely on human assessment, presenting a fundamental challenge.
@@
Large language models achieve empirical success in alignment (Meta, 2024; Yang et al., 2024; DeepSeek-AI, 2024) through Supervised Fine-tuning (SFT) and Reinforcement Learning from Human Feedback (RLHF) techniques. Specifically, SFT (Chung et al., 2022; Wei et al., 2022) finetunes models with human-annotated demonstrations, showing particular efficacy in tasks where humans can effectively showcase desired behaviors. RLHF (Christiano et al., 2023; Ouyang et al., 2022) leverages reinforcement learning with human preference reward models, extending supervision capability to more complex tasks where direct solution generation proves challenging.
@@
However, both approaches rely on human feedback, making them unsustainable for tasks where direct human evaluation becomes infeasible. For example, humans can struggle with the time-consuming tasks, e.g., reviewing extensive content in long-form text (Stiennon et al., 2022) or the expertise required tasks, e.g., to verify solutions in complex mathematical formulations (Li et al., 2024b). Furthermore, as AI capabilities advance beyond human abilities, obtaining reliable supervision signals becomes increasingly challenging, representing the core point of scalable oversight (Casper et al., 2023; Ji et al., 2024; Kenton et al., 2024b).
@@
The principle behind RLHF is that verification is easier than generation (Leike et al., 2018; Irving et al., 2018b). By recognizing critique as a specialized form of generation, we further hypothesize that critique of critique is easier than critique itself. Taking a complex mathematical proof as an example: direct review can be challenging, but assessing its critique is more manageable, as the key steps have already been identified. Moreover, we hypothesize that above difficulty relationship generalizes recursively, where each successive level of meta-evaluation becomes increasingly tractable. This is similar to organizational decision-making procedure, where managers evaluate their subordinates’ assessments rather than directly reviewing complex details. Both above hypotheses, if validated, offer a promising pathway for scalable oversight: while directly evaluating sophisticated AI output may exceed human capabilities, performing higher-order critiques could remain feasible.
@@
To systematically verify and explore these hypotheses, we first conduct Human-Human experiments, where human evaluate human outputs, examining the progression from response to critique and critique-of-critique ($C^{2}$). By comparing accuracy with similar effort, completion time, and confidence levels, we find that higher-order critique contributes to more effective evaluation than direct assessment. Furthermore, we demonstrate the recursive hypothesis by extending experimental settings to deeper critique chains, i.e., critique of critique of critique ($C^{3}$). Inspired by above human-human findings, we further investigate its applicability for supervising AI: when AI generate self-recursive critique, could humans provide effective oversight by evaluating these critique chains? To answer this question, we conduct Human-AI experiments, where human evaluate AI outputs, on tasks where AI outperforms average human. The result is promising and exhibited with models of varying capabilities. Finally, we examine whether AI could achieve effective oversight through recursive self-critiques in AI-AI experiments across models of different capabilities. Our results demonstrate that recursive self-critiquing shows effectiveness in weak-to-strong scenarios, while the optimal critique strategy depends on relative model strengths between the supervised and critic models.
@@
footnotetext: The main idea for this work came from a late-night, insightful discussion between Jie Lou and Xing Yu, which was part of some truly wonderful days.
Figure 1: Overview of the recursive critique framework. Starting from response generation for a given question, each subsequent level performs pair-wise evaluation of outputs from the previous level, forming a recursive criticism chain. $C^{1}$ denotes Critique, $C^{2}$ denotes Critique of Critique, $C^{3}$ denotes Critique of Critique of Critique.
@@ 2 Recursive Self-Critiquing
In this section, we introduce the interaction protocols that define how recursive self-critiquing progresses through multiple levels, from initial response to higher-order critiques. We then present two baselines, majority voting and naive voting, for fair comparison of the recursive critique’s effectiveness.
@@ 2.1 Protocols
As shown in the Figure 1, the hierarchical criticism architecture progresses through multiple levels: from initial response, through first-order critique, to second-order critique of critique ($C^{2}$) and higher-order critiques.
@@
Response represents the initial attempt to answer the question, serving as the foundation of the criticism chain. Each response comprises a complete solution process and its corresponding answer, which is formally defined as:
@@
where $Q$ denotes the input question, $T$ represents the solution process, and $A$ is the final answer. The inclusion of the full solution process enables critiques to evaluate the complete reasoning path rather than merely the result.
@@
The critique evaluates pairs of candidate responses for a given input question, conducting comparative analysis and providing reasoning judgment as follows:
@@
where $R_{1}$ and $R_{2}$ denote candidate responses, $T^{1}$ represents the critique rationale, and $A^{1}$ is the answer to the question based on the responses and critique rationale.
@@
The second-order critique evaluates pairs of critiques, extending the evaluation to a higher level of abstraction, which is defined as:
$C^{2}(Q,R_{1},R_{2},C_{1},C_{2})\rightarrow(T^{2},A^{2})$ (3)
where $C_{1}$ and $C_{2}$ are two critiques of the original responses. $T^{2}$ represents the second-order critique’s analysis of the critiques’ reasoning, and $A^{2}$ denotes the final judgment.
@@
The $n$-th order critique extends this evaluation pattern to higher levels of abstraction, analyzing the assessments from all previous levels, and it is defined as:
$C^{n}(Q,R_{1},R_{2},...,C^{n-1}_{1},C^{n-1}_{2})\rightarrow(T^{n},A^{n})$ (4)
where $C^{n-1}_{1}$ and $C^{n-1}_{2}$ are evaluations from the $(n-1)$-th order critiques, $T^{n}$ represents the $n$-th order critic’s analysis of the previous critiques, and $A^{n}$ denotes the final judgment.
@@ 2.2 Baselines
We introduce two baseline strategies for the rigorous comparison for the recursive critique: majority voting that ensures fair comparison under equivalent computational effort, and naive voting that simply aggregates all available judgments to verify whether recursive critique generates meaningful new insights beyond naive consensus.
@@
Since higher-order critiques are based on all previous evaluation results, we need to consider computational cost for fair comparison. To validate that improvements stem from the recursive structure rather than increased computation, we compare higher-order critiques with majority voting of lower-order evaluations under similar effort. Specifically, let $\epsilon(\cdot)$ denote the computational overhead for each single evaluation. As presented in Figure 1, the total efforts $E(\cdot)$ for different order recursive critique $C^{1}$, $C^{2}$ and $C^{3}$ are defined as:
$E(C^{1})$ $=\epsilon(R_{1})+\epsilon(R_{2})+\epsilon(C^{1})\approx 3\epsilon(R)$ (5)
$E(C^{2})$ $=\epsilon(R_{1})+\epsilon(R_{2})+\epsilon(C^{1}_{1})+\epsilon(C^{1}_{2})+\epsilon(C^{2})\approx 5\epsilon(R)$
$E(C^{3})$ $=\epsilon(R_{1})+\epsilon(R_{2})+\epsilon(C^{1}_{1})+\epsilon(C^{1}_{2})+\epsilon(C^{2}_{1})+\epsilon(C^{2}_{2})+\epsilon(C^{3})\approx 7\epsilon(R)$
Then, we define majority voting. For level $l$, given a set of $n$ evaluations, the majority voting result is:
[…]
where $A^{l}_{i}$ represents the judgment from the $i$-th evaluation at level $l$, and $\mathbbm{1}(\cdot)$ is the indicator function. Intuitively, this formula counts the occurrences of each possible answer among the $n$ evaluations and selects the most frequent one as the final result. In case of ties where multiple answers have the same highest frequency, one is randomly selected. To ensure effort equivalence when comparing with recursive critique at level $l$, we calculate $\text{Major}_{n}^{k}$ where $k<l$ and $n=E(C^{l})/E(C^{k})$. As an example, for a fair comparison, $C^{3}$ should be compared with $\text{Major}_{3}^{2}$ (majority voting among three $C^{2}$ critiques) and $\text{Major}_{5}^{1}$ (majority voting among five $C^{1}$ critiques).
@@ 3.1 Tasks
We select five representative tasks, which call for diverse cognitive capabilities, and remain in moderate difficulty. All tasks include 64 multiple-choice questions.
@@
College English Test Band 6 (CET-6) is a standardized English proficiency assessment for Chinese university students. We select one question per passage from its Careful Reading section; each passage is 400-450 words and includes multiple-choice questions testing main idea comprehension, vocabulary understanding, or inference abilities. Given that few annotators have passed CET-6, these questions present a substantial challenge for them.
@@
The chinese reading comprehension questions are drawn from China’s National College Entrance Examination (Gaokao). These questions evaluate comprehensive reading abilities through textual analysis, logical inference, and meaning interpretation. As our annotators are college graduates who have taken Gaokao before, these questions present moderate difficulty for them.
@@
The mathematics questions are sourced from standardized high school tests in (Zhang et al., 2023). We select first ten multiple-choice problems with moderate difficulty [^1] Question difficulty increases with problem number., as our annotators left campus for a few years and some have non-science backgrounds.
@@
The questions are from China’s National Civil Service Exam, the annual government recruitment test. These questions cover logical reasoning, language understanding, and number analysis. We exclude knowledge questions to focus on cognitive skills rather than factual recall.
@@
These visual tasks from the Civil Service Examination assess logical abilities through non-verbal reasoning without requiring extra knowledge or cultural context.
@@ 3.2 Setup
We recruit 32 participants with bachelor’s degrees, including 22 with STEM backgrounds and 10 with liberal arts backgrounds. Most participants passed CET-4 level English and achieved a score of approximately 100 (out of 150) in high school math exam. Meawhile, most participants have full-time data annotation experience.
Table 1: Human experiment results across response, critique, and $C^{2}$ stages for five tasks. Bold numbers indicate best performance. Majority Voting@$E5$ represents voting results with computational effort equivalent to 5 times of response. Metrics includes average accuracy, voting accuracy, naive voting, confidence (1-5), and completion time (minutes).
CET 6 Response 49.11 55.80 - 3.074 18.36
Critique 58.13 60.78 49.22 3.253 17.03
$C^{2}$ 60.94 60.94 56.25 3.516 15.82
GAOKAO Math Response 66.29 81.81 - 3.201 14.58
Critique 82.50 86.61 66.41 3.863 14.62
$C^{2}$ 90.62 90.62 81.25 3.979 15.48
GAOKAO Chinese Response 71.56 79.69 - 3.822 17.81
Critique 78.65 84.38 64.84 4.026 13.91
$C^{2}$ 84.38 84.38 77.34 4.078 10.25
Figure Reasoning Response 65.00 78.12 - 3.888 16.74
Critique 75.00 77.08 65.62 4.213 16.01
$C^{2}$ 79.69 79.69 72.66 4.313 15.02
KAOGONG Response 69.69 83.59 - 3.828 16.26
critique 84.38 84.90 70.31 4.031 15.48
$C^{2}$ 85.94 85.94 82.81 4.031 12.58
Table 2: Human experiment results across response, critique, $C^{2}$, and $C^{3}$ stages for two tasks. Bold numbers indicate best performance. Majority Voting@$E7$ represents voting results with computational effort equivalent to 7 times of response. Metrics includes accuracy, voting accuracy, naive voting, confidence (1-5), and completion time (minutes).

### v4

@@ Abstract
As AI capabilities increasingly surpass human proficiency in complex tasks, current alignment techniques, including SFT and RLHF, face fundamental challenges in ensuring reliable oversight. These methods rely on direct human assessment and become impractical when AI outputs exceed human cognitive thresholds. In response to this challenge, we explore two hypotheses: (1) Critique of critique can be easier than critique itself, extending the widely-accepted observation that verification is easier than generation to the critique domain, as critique itself is a specialized form of generation; (2) This difficulty relationship holds recursively, suggesting that when direct evaluation is infeasible, performing higher-order critiques (e.g., critique of critique of critique) offers a more tractable supervision pathway. We conduct Human-Human, Human-AI, and AI-AI experiments to investigate the potential of recursive self-critiquing for AI supervision. Our results highlight recursive critique as a promising approach for scalable AI oversight.
Machine Learning, ICML
@@ 1 Introduction
Supervision signals are fundamental to AI alignment (Bowman et al., 2022), providing the ground truth or preference data necessary to train models that behave in accordance with human expectations. The nature and accessibility of these supervision signals, however, vary substantially across different application domains. From a supervision acquisition perspective, tasks can be categorized into two types: (1) tasks with well-defined criteria, where ground truth can be deterministically obtained with low computational overhead, e.g., Go games and mathematical problems (Silver et al., 2017; Lightman et al., 2023); (2) tasks involving subjectivity or complex evaluation frameworks, such as business strategy and product design (Ouyang et al., 2022). The latter type is more prevalent in real-world applications and predominantly relies on human assessment, presenting a fundamental challenge.
@@
Current alignment techniques, particularly Supervised Fine-tuning (SFT) and Reinforcement Learning from Human Feedback (RLHF), have achieved empirical success with large language models (Meta, 2024; Yang et al., 2024; DeepSeek-AI, 2024). SFT (Chung et al., 2022; Wei et al., 2022) finetunes models with human-annotated demonstrations, showing particular efficacy in tasks where humans can effectively showcase desired behaviors. RLHF (Christiano et al., 2023; Ouyang et al., 2022) employs reinforcement learning with human preference reward models based on pairwise comparisons, extending supervision to more complex tasks where direct solution generation is challenging.
@@
However, both approaches rely on direct human feedback, making them unsustainable for tasks where human evaluation becomes infeasible. For example, humans can struggle with time-consuming tasks such as reviewing extensive long-form text (Stiennon et al., 2022)r expertise-intensive tasks such as verifying solutions to complex mathematical problems (Li et al., 2024b). Furthermore, as AI capabilities advance beyond human abilities, obtaining reliable supervision signals becomes increasingly challenging, representing the central problem of scalable oversight (Casper et al., 2023; Ji et al., 2024; Kenton et al., 2024b).
@@
The underlying insight of RLHF is that verification is easier than generation (Leike et al., 2018; Irving et al., 2018b). By recognizing critique as a specialized form of generation, we further hypothesize that critique of critique is easier than critique itself. Taking a complex mathematical proof as an example: while direct review can be challenging, assessing its critique is more manageable, as the key steps have already been identified. Moreover, we hypothesize that this difficulty relationship generalizes recursively, where each successive level of meta-evaluation becomes increasingly tractable. This resembles organizational decision-making processes, where managers evaluate their subordinates’ assessments rather than directly reviewing complex details. These hypotheses, if validated, offer a promising pathway for scalable oversight: while directly evaluating sophisticated AI output may exceed human capabilities, performing higher-order critiques could remain feasible.
@@
To systematically verify these hypotheses, we first conduct Human-Human experiments where humans evaluate human outputs. We examine the progression from response to critique and then to critique-of-critique ($C^{2}$). By comparing accuracy under similar computational effort, completion time, and confidence levels, we find that higher-order critiques contribute to more effective evaluation than direct assessment. Furthermore, we demonstrate the recursive nature of this relationship by extending experiments to deeper critique chains, i.e., critique of critique of critique ($C^{3}$). Inspired by these human-human findings, we further investigate their applicability for supervising AI: when AI generates self-recursive critiques, can humans provide effective oversight by evaluating these critique chains? To answer this question, we conduct Human-AI experiments, where humans evaluate AI outputs on tasks where AI outperforms average humans. The results are promising across models of varying capabilities. Finally, we examine whether AI can achieve effective oversight through recursive self-critiques in AI-AI experiments across models of different capabilities. Our results demonstrate that recursive self-critiquing is effective in weak-to-strong scenarios, while the optimal critique strategy depends on the relative capabilities between supervised and critic models.
@@ 2 Recursive Self-Critiquing
In this section, we introduce the protocols for recursive self-critiquing across multiple evaluation levels, spanning initial response through higher-order critiques. We then present majority voting and naive voting as two representative baselines to provide fair comparisons for evaluating the effectiveness of recursive critique.
Figure 1: Overview of the recursive critique framework. Starting from response generation for a given question, each subsequent level performs pair-wise evaluation of outputs from the previous level, forming a recursive critique chain. $C^{1}$ denotes Critique, $C^{2}$ denotes Critique of Critique, $C^{3}$ denotes Critique of Critique of Critique.
@@ 2.1 Protocols
As shown in Figure 1, the hierarchical criticism architecture progresses through multiple levels: from initial response, through first-order critique, to second-order critique of critique ($C^{2}$) and higher-order critiques. Our protocols follow standard RLHF practice (Ouyang et al., 2022), employing pairwise comparisons at each critique level. This approach leverages humans’ cognitive advantage in relative assessment over absolute evaluation (Jones and Inglis, 2015; Kelly et al., 2022), making recursive evaluation more tractable at each level. Moreover, this design facilitates consistency between human and AI experiments, as the latter requires pairwise preference data for reward model training.
@@
Response is the initial attempt to answer the question, serving as the foundation of the critique chain. Each response comprises a complete solution process and its corresponding answer:
@@
where $Q$ denotes the input question, $T^{0}$ represents the solution process which may include reasoning steps, justifications, and intermediate calculations, and $A^{0}$ is the final answer. Including the full solution process rather than merely the final answer enables critiques to better assess the correctness of each response by examining logical consistency, key step validity, and other aspects of the solution.
@@
The first-order critique evaluates pairs of candidate responses for a given input question, conducting comparative analysis and providing reasoned judgment:
@@
where $R_{1}$ and $R_{2}$ denote two candidate responses, $T^{1}$ represents the critique rationale explaining which response is better and why, and $A^{1}$ is the final answer determined based on the critique analysis.
@@
The second-order critique evaluates pairs of first-order critiques, extending the evaluation to a higher level of abstraction:
$C^{2}(Q,R_{1},R_{2},C^{1}_{1},C^{1}_{2})\rightarrow(T^{2},A^{2})$ (3)
where $C^{1}_{1}$ and $C^{1}_{2}$ are two first-order critiques of the original responses, $T^{2}$ represents the analysis comparing the quality and validity of these critiques, and $A^{2}$ denotes the final answer determined by the superior critique.
@@
The $n$-th order critique continues this recursive process, leveraging assessments from all previous levels for evaluating pairs of (n1)(n-1)-th order critiques and reaching conclusions at this level:
$C^{n}(Q,R_{1},R_{2},C^{1}_{1},C^{1}_{2},\ldots,C^{n-1}_{1},C^{n-1}_{2})\rightarrow(T^{n},A^{n})$ (4)
where $C^{n-1}_{1}$ and $C^{n-1}_{2}$ are two (n1)(n-1)-th order critiques, $T^{n}$ represents the analysis comparing these critiques, and $A^{n}$ denotes the final answer derived from this comprehensive evaluation.
@@ 2.2 Baselines
We introduce two representative baseline strategies for rigorous comparison with recursive critique. The first is majority voting, which selects the most frequent answer from multiple evaluations. This baseline ensures fair comparison under equivalent computational effort. The second is naive voting, which performs direct aggregation of all available judgments from previous stages. This approach verifies whether recursive critique generates meaningful insights beyond simple consensus.
@@
Since higher-order critiques are based on lower-order evaluation results, direct comparison between them would be unfair due to differing computational costs. To verify that the recursive structure achieves performance improvements by reducing supervision difficulty rather than merely benefiting from increased computational effort, we compare higher-order critiques with lower-order critiques under approximately equivalent computational effort. We achieve this through majority voting baselines (Wang et al., 2023) that aggregate multiple lower-order evaluations to match the computational cost of higher-order critiques.
Specifically, let $\epsilon(\cdot)$ denote the computational overhead for each evaluation. In AI experiments, this typically represents inference cost, while in human experiments, it’s more closely captured by annotation time spent on each evaluation task. As presented in Figure 1, the total computational effort $E(\cdot)$ for different-order recursive critiques $C^{1}$, $C^{2}$, and $C^{3}$ can be estimated as:
$E(C^{1})$ $=2\epsilon(R)+\epsilon(C^{1})\approx 3\epsilon(R)$ (5)
$E(C^{2})$ $=2\epsilon(R)+2\epsilon(C^{1})+\epsilon(C^{2})\approx 5\epsilon(R)$
$E(C^{3})$ $=2\epsilon(R)+2\epsilon(C^{1})+2\epsilon(C^{2})+\epsilon(C^{3})\approx 7\epsilon(R)$
We then define majority voting. For level $l$, given a set of $n$ evaluations, the majority voting result is:
$\text{Major}_{n}^{l}(\mathcal{A})=\underset{a}{\operatorname{argmax}}\sum_{i=1}^{n}\mathbbm{1}(A^{l}_{i}=a)$ (6)
where $A^{l}_{i}$ represents the judgment from the $i$-th evaluation at level $l$, and $\mathbbm{1}(\cdot)$ is the indicator function. This formula counts the occurrences of each possible answer among the $n$ evaluations and selects the most frequent one as the final result. In case of ties, where multiple answers have the same highest frequency, one is randomly selected. To enable fair comparison with recursive critique at level $l$ under approximately equivalent effort, we calculate $\text{Major}_{n}^{k}$ where $k<l$ and $n=E(C^{l})/E(C^{k})$. For example, $C^{3}$ should be compared with $\text{Major}_{3}^{2}$ (majority voting among three $C^{2}$ critiques) and $\text{Major}_{5}^{1}$ (majority voting among five $C^{1}$ critiques). Critically, majority voting aggregates independent evaluations without the structured pairwise comparison that defines recursive critique, allowing us to isolate whether improvements stem from the recursive structure versus computational scaling.
@@ 3 Is Recursive Critique Increasingly Easier?
Table 1: Human experiment results across response, critique, and $C^{2}$ stages for five tasks. Bold numbers indicate best performance. Majority Voting@$E5$ represents voting results with computational effort equivalent to 5 times of response. Metrics include average accuracy, voting accuracy, naive voting, confidence (1-5), and completion time (minutes).
CET-6 Response 49.11 55.80 3.074 18.36
Critique 58.13 60.78 49.22 3.253 17.03
$C^{2}$ 60.94 56.25 3.516 15.82
GAOKAO Math Response 66.29 81.81 3.201 14.58
Critique 82.50 86.61 66.41 3.863 14.62
$C^{2}$ 90.62 81.25 3.979 15.48
GAOKAO Chinese Response 71.56 79.69 3.822 17.81
Critique 78.65 84.38 64.84 4.026 13.91
$C^{2}$ 84.38 77.34 4.078 10.25
Figure Reasoning Response 65.00 78.12 3.888 16.74
Critique 75.00 77.08 65.62 4.213 16.01
$C^{2}$ 79.69 72.66 4.313 15.02
KAOGONG Response 69.69 83.59 3.828 16.26
Critique 84.38 84.90 70.31 4.031 15.48
$C^{2}$ 85.94 82.81 4.031 12.58
@@ 3.1 Tasks
We select five representative tasks that require diverse cognitive capabilities while maintaining moderate difficulty. These tasks span multiple domains, including language comprehension, mathematical reasoning, logical analysis, and visual reasoning, to test the generalizability of recursive critique framework across different cognitive skills. All tasks include 64 multiple-choice questions. Each task consists of 64 multiple-choice questions.
@@
The College English Test Band 6 (CET-6) is a standardized English proficiency assessment for Chinese university students. From its Careful Reading section, we select one question per passage; each passage contains 400-450 words and includes multiple-choice questions that test main idea comprehension, vocabulary understanding, and inference abilities. This task requires English language proficiency, reading comprehension skills, and analytical reasoning to extract meaning from complex texts. As few of our annotators have passed CET-6, these questions present substantial challenges.
@@
The Chinese reading comprehension questions are drawn from China’s National College Entrance Examination (Gaokao). These questions demand accurate comprehension of the original text and logical reasoning capabilities for answer selection. Since our annotators are college graduates who previously took the Gaokao, these questions present moderate difficulty.
@@
The mathematics questions are sourced from standardized high school tests (Zhang et al., 2023). Since problem difficulty typically increases with question number and considering that our annotators graduated several years ago with some having non-science backgrounds, we select the first ten multiple-choice problems to ensure moderate difficulty for them. These questions require mastery of mathematical concepts and formulas as well as the ability to apply mathematical reasoning to solve problems.
@@
The questions are sourced from China’s National Civil Service Exam, the annual government recruitment test. These questions assess logical reasoning, language understanding, and numerical analysis skills. We exclude knowledge-based questions to focus on cognitive abilities requiring analytical thinking and problem-solving rather than factual recall.
@@
These visual tasks from the Civil Service Examination assess logical abilities through non-verbal reasoning without requiring domain-specific knowledge or cultural context, demanding spatial reasoning skills, pattern recognition, and abstract thinking capabilities.
@@ 3.2 Setup
We recruit 32 participants with bachelor’s degrees: 22 from STEM backgrounds and 10 from liberal arts backgrounds. Most participants passed CET-4 and scored approximately 100 out of 150 on their high school mathematics exams. All participants have prior experience in data annotation and are employed full-time for this study.
Execution
We develop standardized guidelines for all tasks using instructions and examples, detailed in Appendix A. Tasks are organized into data packages with specified submission deadlines, and annotators are randomly assigned across different critique levels to ensure participation in all stages. To maintain efficiency, we set a 20-minute time limit for each question at every stage, managed through flexible package-level deadlines that allow annotators to allocate time as needed. Annotators complete a predetermined number of tasks daily within their scheduled working hours. We conduct regular feedback sessions to gather suggestions for refining procedures and guidelines, and assign dedicated staff to oversee process management and quality control.
Metrics
We assess the effectiveness of recursive critique through three metrics: (1) accuracy measures consistency with ground truth answers; (2) completion time records the duration of the entire evaluation process; (3) confidence reflects participants’ self-assessed certainty in their final answers on a five-point scale.
3.3 Critique of Critique can be Easier than Critique
We validate the hypothesis that critique of critique is easier than critique across five tasks. The results in Table 1 show consistent improvements from response to critique to $C^{2}$ stages. Taking GAOKAO Math as an example, average accuracy improves from 66.29% (response) to 82.50% (critique) and further to 90.62% ($C^{2}$), while completion time remains stable or slightly decreases (e.g., from 18.36 to 15.82 minutes for CET-6). Under comparable effort, majority voting shows similar trends. For instance, accuracy improves from 81.81% (response) through 86.61% (critique) to 90.62% ($C^{2}$) in GAOKAO Math, demonstrating the advantage of higher-order critique. Compared to naive voting, average accuracy consistently outperforms. Taking GAOKAO Math as an example, naive voting achieves only 66.41% at the critique stage and 81.25% at $C^{2}$, significantly lower than the average accuracy of 90.62%. These results validate that recursive critique generates new insights rather than merely aggregating previous judgments. Moreover, annotator confidence shows steady improvement across stages, suggesting that higher-order critique becomes more tractable.
3.4 Recursive Critique Remains Consistently Easier
We extend recursive critique to third-order critique ($C^{3}$) on two representative tasks. As shown in Table 2, accuracy improves continuously at the $C^{3}$ level in both tasks, with CET-6 increasing from 60.94% at $C^{2}$ to 67.19%, and GAOKAO Math from 90.62% to 93.75%. Under comparable computational effort, majority voting shows similar improvements, reaching 67.19% for CET-6 and 93.75% for GAOKAO Math. However, naive voting achieves substantially lower performance than the average accuracy at each critique level. Meanwhile, confidence scores improve while completion time decreases across critique levels. These results demonstrate that recursive critique remains consistently easier and that its benefits extend beyond mere computational scaling or consensus aggregation.
Table 2: Human experiment results across response, critique, $C^{2}$, and $C^{3}$ stages for two tasks. Bold numbers indicate best performance. Majority Voting@$E7$ represents voting results with computational effort equivalent to 7 times of response. Metrics include accuracy, majority voting accuracy, naive voting, confidence (1-5), and completion time (minutes).